### v2

@@ I Introduction
In plasma physics, transport induced by electrostatic waves is most often modeled as a quasilinear diffusion [12]. Although quasilinear theory has been the subject of considerable debate [13], several papers provided a detailed discussion of its relevance for a nearly stationary spectrum (see for example Refs. [1, 5, 14]). In particular, Ref. 1 shows that two types of diffusion may exist. In the first one, which is quasilinear and closer to the situation considered in the initial theory [12], the particle motion may be described by making use of a first order perturbation analysis. It may be followed by a second, so-called “chaotic” diffusion regime, when the particle motion looks erratic. In this second regime, the diffusion coefficient may be different from the quasilinear one, although usually of the same order. Moreover, as shown in Refs. [1, 3], for large enough wave amplitudes, the initial quasilinear regime persists until the onset of the second, chaotic regime. In this situation, quasilinear diffusion is valid until the particules reach the phase space boundaries. Now, all the aforementioned results were derived for the dynamics defined by the following Hamiltonian, assumed to be paradigmatic,
@@
where the $\varphi_{n}$s were random initial phases. The key point of the derivation was the proof of locality in phase velocity for the wave-particle interaction [15]. Here, locality means that, at all times, transport is essentially due to those waves whose phase velocity differ by no more than $\Delta v_{R}\approx 6A^{2/3}$ from the instantaneous particle speed. Note that the quasilinear diffusion coefficient, as derived in Ref. 12 for a continuous spectrum, accounts only for waves with phase velocity exactly matching the particle speed, thus representing an infinitely local interaction. This would not make sense for a discrete spectrum and, for such a wave spectrum, the expression of the quasilinear diffusion coefficient was derived in Ref. 18 and recalled in Section III, Eq. (39).
@@
    Does there exist a broad class of Hamiltonian dynamics whose transport properties are exactly the same as those derived from $H_{p}$ ?
@@
    Talking about diffusion is somewhat abusive for a finite discrete wave spectrum because particles only have access to a limited region of phase space, which raises the following question : can one derive an analytical expression for the time evolution of the velocity distribution function that accounts for phase space boundaries ?
To the best of the author’s knowledge, no previous article specifically addressed the first fourth and sixth issues. In Ref. 5, Vlasov simulations showed the relevance of quasilinear theory to model the warm beam-plasma instability in the strongly nonlinear regime, while non-quasilinear evolution was observed for intermediate nonlinear regimes, thus supporting some of the results derived from the Hamiltonian $H_{p}$. Moreover, on a more theoretical perspective, several papers, e.g. Ref. 14, addressed the relevance of the initial quasilinear diffusion regime for more general wave spectra than that considered in Hamiltonian $H_{p}$. However, the authors of these papers implicitly made hypotheses that led to the validity of the initial quasilinear regime. In this article, we aim to address the 6 aforementioned issues for the general class of wave spectra leading to the dynamics defined by the following Hamiltonian,
@@
where the $\varphi_{n}$s remain random initial phases, while the $a_{n}$s, $k_{n}$s and $\omega_{n}$s may be chosen arbitrarily. As for the pre-factor, $A$, it is introduced as a normalization constant to allow direct comparison with the results derived from Hamiltonian $H_{p}$. For the general Hamiltonian $H$, we theoretically establish the conditions on the relative variations of the wave amplitudes, wave numbers, and wave frequencies required for locality to hold. We further show that the presence of an initial quasilinear diffusion regime is closely linked to the validity of locality. In addition, we derive the criteria under which the quasilinear diffusion coefficient for $H$ coincides with that for $H_{p}$, namely $D_{\rm{QL}}=\pi A^{2}/2$, and under which the overlap parameter, $s=2\sqrt{A}\left(\sqrt{a_{n}}+\sqrt{a_{n+1}}\right)/\delta v_{\phi_{n}}$, $\delta v_{\phi_{n}}=|d_{n}(\omega_{n}/k_{n})|,$ also matches that of $H_{p}$, i.e., $s=4\sqrt{A}$. When these conditions are met, the transport properties for Hamiltonian $H$ should be very similar to those of Hamiltonian $H_{p}$. In particular, they should be uniform in phase space [at least in the region of phase space corresponding to $v\lesssim\max(\omega_{n}/k_{n})$ and $v\gtrsim\min(\omega_{n}/k_{n})$]. Consequently, if a Fokker-Planck equation is valid, it reduces to a simple diffusion equation, $\partial_{t}f=D\partial^{2}_{v}f$, where $D$ is independent of $v$.
@@
Now, diffusion cannot last for all times because particles eventually hit the boundaries of phase space, as illustrated in Fig. 3. These boundaries are the so-called KAM tori [19] which exist for $v\gtrsim\max(\omega_{n}/k_{n})$ and $v\lesssim\min(\omega_{n}/k_{n})$ and which cannot be crossed by any orbit. Consequently, one cannot simply derive the velocity distribution function by solving the diffusion equation, $\partial_{t}f=D\partial^{2}_{v}f$. In this paper, we provide a heuristic expression for the time variation of the distribution function, which accounts for phase space boundaries, and whose accuracy is tested against results from test particle simulations.
@@ II Bounded chaotic transport
In this section, we qualitatively discuss the physical effect we aim to quantify : a bounded chaotic transport in velocity. The discussion is based on Poincaré sections for the dynamics ruled by the Hamiltonian $H_{p}$, Eq. (1). This amounts to plotting the phase space coordinates $\left[x\left(t_{n}\right),v\left(t_{n}\right)\right]$ at discrete times $t_{n}=2\pi n$. The positions $x(t_{n})$ are plotted modulo $2\pi$ so that $0\leq x(t_{n})\leq 2\pi$.
@@
When $N>1$, if the phase portrait were a mere superposition of $N$ pendulum-like phase portraits, it would consist of $N$ sets of trapped and untrapped orbits, separated by $N$ separatrices, centered about $v=v_{\phi_{n}}$ and of width $4\sqrt{Aa_{n}}$. Clearly, this picture has no chance to be correct when the separatrices overlap, that is, when the Chirikov overlap parameter, $s$, is larger than unity, where
$s=2\sqrt{A}\frac{\sqrt{a_{n}}+\sqrt{a_{n+1}}}{\delta v_{\phi_{n}}},$ (3)
with $\delta v_{\phi_{n}}=|d_{n}(\omega_{n}/k_{n})|$. For the Hamiltonian $H_{p}$, $s=4\sqrt{A}$.
@@
The first and third initiation conditions lie, respectively, within the separatrices associated with $n=11$ and $n=12$, so that the corresponding orbits are expected to be the counterparts of trapped orbits for the pendulum. The second initial condition is located above the separatrix for $n=11$, so that the corresponding orbit is expected to be the counterpart of an untrapped orbit for the pendulum. The fourth and fifth initial conditions are situated, respectively, at the X-points of the separatrices corresponding to $n=11$ and $n=12$. As for the sixth and seventh initial conditions, depending on the value of $A$, they may lie either within or above the separatrix corresponding to $n=11$.
@@ III.1 Validity of locality
In this Paragraph, we discuss the relevance of locality for the Hamiltonian $H$ defined by Eq. (2). To do so, we use the same procedure as in Ref. 15, i.e., we make use of a perturbation analysis about a given velocity, $v_{0}$. This lets us conclude whether there exists a range in velocity, $\Delta v_{R}$, such that the waves with phase velocity $v_{\phi}$ satisfying $|v_{0}-v_{\phi}|>\Delta v_{R}$ have, on the average, a negligible impact on transport. Before proceeding to the analysis, we renormalize velocities to the, still unknown, range in velocity, $\Delta v_{R}$. Hence, we define the following new variables,
@@
Now, to prove locality, we need to prove that there exists a canonical change of variables $(x_{1},v_{1})\to(x^{\prime},v^{\prime})$, defined from a generative function, $\Phi(x_{1},v^{\prime},t)\equiv x_{1}v^{\prime}+\sum_{l}\varepsilon^{l}\Phi_{l}$ (the small parameter of the expansion, $\varepsilon$, being still to define) such that, in the new variables,
@@
where the $\mathcal{A}_{n}$s, $\kappa_{n}$s, $\varpi_{n}$s and $\psi_{n}$s are amplitudes, wave numbers, frequencies and phases which are derived from the perturbative expansion, and where there exists an upper bound for the remainder, $R_{l}$, independent of the number of modes, $N$. For the Hamiltonian $H_{p}$, Eq. (1), considered in Ref. 15, the terms of the perturbative expansion could be bounded by those of a Gevrey series, which allowed to provide an exponentially small estimate for $\varepsilon^{l+1}R_{l+1}$, with $\varepsilon=A/\Delta v_{R}^{3/2}$.
@@
Note that $\Delta_{n_{i}}>1$ when $v^{\prime}=v_{0}/\Delta v_{R}$. Since we are only interested in statistical properties then, just like in Ref. 15, we estimate the remainder by its mean value with regards to the phases $\varphi_{n}$. When $v^{\prime}\approx v_{0}/\Delta v_{R}$,
@@
where $x_{\min}=\min(l)$ and $x_{\max}=\max(l)$ for each $l$ fulfilling Eq. (25). Clearly, the property of locality is satisfied if one can find an upper bound for the remainder which is independent of the extent, in phase velocity, of the wave spectrum. This is true if the integral in the right hand side of Eq. (26) converges when $x_{\max}\to\infty$. This happens only when $A^{2}(x)/x$ goes to zero when $x$ goes to infinity, i.e., $A^{2}(v_{\phi})/v_{\phi}$ goes to zero when $v_{\phi}$ goes to infinity. Hence, the property of locality is not always true, depending on the wave spectrum.
@@
$\Delta v_{R}\sim\left(\overline{A^{2}_{v}}\right)^{1/3},$ (28)
which lets naturally choose $\varepsilon=\sqrt{\overline{A^{2}_{v}}}/\Delta v_{R}^{3/2}$ for the small parameter of the perturbation analysis. This is a plain generalization of the small parameter introduced in Ref. 15 for the Hamiltonian $H_{p}$.
@@ III.2.1 The quasilinear diffusion coefficient for a discrete wave spectrum
$\Delta v_{R}\sim\left(D_{\rm{QL}}/\overline{k}\right)^{1/3},$ (36)
and is reminiscent of the results previously obtained by Dupree [16], although derived in a completely different sprit. Now,
@@
$D_{\rm{QL}}=\frac{\pi}{2}\left|\frac{k_{n}A^{2}a_{n}^{2}}{\delta v_{\phi_{n}}}\right|_{v_{\phi_{n}}\approx v}.$ (39)
@@
which is the same formula as that already published in Ref. 18.
@@ III.2.2 Validity of quasilinear diffusion and connection with locality
In order to derive quasilinear diffusion in Paragraph III.2.1, we had to assume that the dominant contribution to $\langle\Delta v^{2}\rangle$ was from the terms $n$ such that $v_{\phi_{n}}\approx v_{0}$, and that these terms all had about the same amplitude. This may only be true if the series $\sum_{n}\left[k_{n}a_{n}/(k_{n}v_{0}-\omega_{n})\right]^{2}$ converges. This is exactly the same condition as Eq. (41) for locality. Consequently, the initial quasilinear regime of diffusion is only expected when wave-particle interaction is local.
@@ IV.1.1 $A=1$, $v_{0}=50$
Figure 4: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=-1$, $A=1$ and $v_{0}=50$. The blue solid line is for $N=142$, the green dashed line for $N=448$, the red dashed line for $N=1415$ and the black dashed-dotted line for $N=14143$.
@@ IV.2.1 $A=1$, $v_{0}=50$
Figure 7: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=-2$, $A=1$ and $v_{0}=50$. The red dashed solid line is for $N=165$, the black solid line for $N=1041$.
@@
Figure 8: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=-2$, $A=11.5$ and $v_{0}=200$. The red dashed solid line is for $N=660$, the black solid line for $N=4163$.

### v3

@@ I Introduction
In plasma physics, transport induced by electrostatic waves is most often modeled as a quasilinear diffusion [12]. Although quasilinear theory has been the subject of considerable debate [13], several papers provided a detailed discussion of its relevance for a nearly stationary spectrum (see for example Refs. [1, 5, 14, 15]). In particular, Ref. 1 shows that two types of diffusion may exist. In the first one, which is quasilinear and closer to the situation considered in the initial theory [12], the particle motion may be described by making use of a first order perturbation analysis. It may be followed by a second, so-called “chaotic” diffusion regime, when the particle motion looks erratic. In this second regime, the diffusion coefficient may be different from the quasilinear one, although usually of the same order. Moreover, as shown in Refs. [1, 3], for large enough wave amplitudes, the initial quasilinear regime persists until the onset of the second, chaotic regime. In this situation, quasilinear diffusion is valid until the particles orbits reach the phase space boundaries. Now, all the aforementioned results were derived for the dynamics defined by the following Hamiltonian, assumed to be paradigmatic,
@@
where the $\varphi_{n}$s were random initial phases. The key point of the derivation was the proof of locality in phase velocity for the wave-particle interaction [16]. Here, locality means that, at all times, transport is essentially due to those waves whose phase velocity differ by no more than $\Delta v_{R}\approx 6A^{2/3}$ from the instantaneous particle speed. Note that the quasilinear diffusion coefficient, as derived in Ref. 12 for a continuous spectrum, accounts only for waves with phase velocity exactly matching the particle speed, thus representing an infinitely local interaction. This would not make sense for a discrete spectrum and, for such a wave spectrum, the expression of the quasilinear diffusion coefficient was derived in Ref. 22 and recalled in Section III, Eq. (39).
@@
    Does there exist a broad class of Hamiltonian dynamics whose transport properties are very similar to those derived from $H_{p}$ ?
@@
    Talking about diffusion is somewhat abusive for a finite discrete wave spectrum because particles orbits only have access to a limited region of phase space, which raises the following question : can one derive an analytical expression for the time evolution of the velocity distribution function that accounts for phase space boundaries ?
To the best of the author’s knowledge, no previous article specifically addressed the first, fourth, and sixth issues. In Ref. 5, Vlasov simulations showed the relevance of quasilinear theory to model the warm beam-plasma instability in the strongly nonlinear regime, while non-quasilinear evolution was observed for intermediate nonlinear regimes, thus supporting some of the results derived from the Hamiltonian $H_{p}$. Moreover, on a more theoretical perspective, several papers, e.g. Ref. 14, addressed the relevance of the initial quasilinear diffusion regime for more general wave spectra than that considered in Hamiltonian $H_{p}$. However, the authors of these papers implicitly made hypotheses that led to the validity of the initial quasilinear regime. In this article, we aim to address the 6 aforementioned issues for the general class of wave spectra leading to the dynamics defined by the following Hamiltonian,
@@
where the $\varphi_{n}$s remain random initial phases, while the $a_{n}$s, $k_{n}$s and $\omega_{n}$s may be chosen arbitrarily. As for the pre-factor, $A$, it is introduced as a normalization constant to allow direct comparison with the results derived from Hamiltonian $H_{p}$. For the general Hamiltonian $H$, we theoretically establish the conditions on the relative variations of the wave amplitudes, wave numbers, and wave frequencies required for locality to hold. We further show that the presence of an initial quasilinear diffusion regime is closely linked to the validity of locality. In addition, we derive the criteria under which the quasilinear diffusion coefficient and the Chirikov overlap parameter [17] for $H$ coincide with those for $H_{p}$. When the indices are ordered in such a way that $v_{\phi_{n}}\equiv\omega_{n}/k_{n}$ is an increasing function of $n$, the Chirikov overlap parameter reads
$s(v)=2\sqrt{A}\frac{\sqrt{a_{n^{*}}}+\sqrt{a_{n^{*}+1}}}{\delta v_{\phi_{n^{*}}}},$} (3)
where $\delta v_{\phi_{n^{*}}}=v_{\phi_{n^{*}+1}}-v_{\phi_{n^{*}}}$, and $n^{*}$ is such that $v_{\phi_{n^{*}}}\leq v\leq v_{\phi_{n^{*}+1}}$. Note that for $H_{p}$ and when $\min(v_{\phi_{n}})\lesssim v\lesssim\max(v_{\phi_{n}})$, $s=4\sqrt{A}$ is uniform in $v$, and the quasilinear diffusion coefficient, $D_{\rm{QL}}$ given by Eq. (39), is also uniform in $v$, $D_{\rm{QL}}=\pi A^{2}/2$. When $D_{\rm{QL}}$ and $s$ are the same for $H$ and $H_{p}$, the transport properties for the dynamics defined by these two Hamiltonians should be very similar. In particular, they should be uniform in phase space [at least in the region of phase space corresponding to $\min(\omega_{n}/k_{n})\lesssim v\lesssim\max(\omega_{n}/k_{n})$]. Consequently, if a Fokker-Planck equation is valid, it reduces to a simple diffusion equation, $\partial_{t}f=D\partial^{2}_{v}f$, where $D$ is independent of $v$.
@@
Now, diffusion cannot last for all times because the particles orbits eventually hit the boundaries of phase space, as illustrated in Fig. 3. These boundaries are the so-called KAM tori [18] which exist for $v\gtrsim\max(\omega_{n}/k_{n})$ and $v\lesssim\min(\omega_{n}/k_{n})$ and which cannot be crossed by any orbit. Consequently, one cannot simply derive the velocity distribution function by solving the diffusion equation, $\partial_{t}f=D\partial^{2}_{v}f$. In this paper, we provide a heuristic expression for the time variation of the distribution function, which accounts for phase space boundaries, and whose accuracy is tested against results from test particle simulations.
@@ II Bounded chaotic transport
In this section, we qualitatively discuss the physical effect we aim to quantify : a bounded chaotic transport in velocity. The discussion is based on Poincaré sections for the dynamics ruled by the Hamiltonian $H_{p}$, Eq. (1). This amounts to plotting the phase space coordinates $\left[x\left(t_{n}\right),v\left(t_{n}\right)\right]$ at discrete times $t_{n}=2\pi n$. The positions $x(t_{n})$ are plotted modulo $2\pi$ so that $0\leq x(t_{n})\leq 2\pi$. In Figs. 2 and 3, the phase space coordinates are computed by solving numerically the Hamilton equations for $H_{p}$ using a leapfrog integrator [19] with the time step $\delta t=2\pi\times 10^{-3}$, while the total integration time is $\Delta t=2\pi\times 10^{4}$.
@@
When $N>1$, if the phase portrait were a mere superposition of $N$ pendulum-like phase portraits, it would consist of $N$ sets of trapped and untrapped orbits, separated by $N$ separatrices, centered about $v=v_{\phi_{n}}$ and of width $4\sqrt{Aa_{n}}$. Clearly, this picture has no chance to be correct when the separatrices overlap, that is, when the Chirikov overlap parameter $s$ defined by Eq. (3) is larger than unity. For the Hamiltonian $H_{p}$, $s=4\sqrt{A}$ whenever $1\lesssim v\lesssim N$.
@@
The first and third initial conditions lie, respectively, within the separatrices associated with $n=11$ and $n=12$, so that the corresponding orbits are expected to be the counterparts of trapped orbits for the pendulum. The second initial condition is located above the separatrix for $n=11$, so that the corresponding orbit is expected to be the counterpart of an untrapped orbit for the pendulum. The fourth and fifth initial conditions are situated, respectively, at the X-points of the separatrices corresponding to $n=11$ and $n=12$. As for the sixth and seventh initial conditions, depending on the value of $A$, they may lie either within or above the separatrix corresponding to $n=11$.
@@ III.1 Validity of locality
In this Paragraph, we discuss the relevance of locality for the Hamiltonian $H$ defined by Eq. (2). To do so, we use the same procedure as in Ref. 16, i.e., we make use of a perturbation analysis about a given velocity, $v_{0}$. This lets us conclude whether there exists a range in velocity, $\Delta v_{R}$, such that the waves with phase velocity $v_{\phi}$ satisfying $|v_{0}-v_{\phi}|>\Delta v_{R}$ have, on the average, a negligible impact on transport. Before proceeding to the analysis, we normalize velocities to the, still unknown, range in velocity, $\Delta v_{R}$. Hence, we define the following new variables,
@@
Now, to prove locality, we need to prove that there exists a canonical change of variables $(x_{1},v_{1})\to(x^{\prime},v^{\prime})$, defined from a generating function, $\Phi(x_{1},v^{\prime},t)\equiv x_{1}v^{\prime}+\sum_{l}\varepsilon^{l}\Phi_{l}$ (the small parameter of the expansion, $\varepsilon$, being still to define) such that, in the new variables,
@@
where the $\mathcal{A}_{n}$s, $\kappa_{n}$s, $\varpi_{n}$s and $\psi_{n}$s are amplitudes, wave numbers, frequencies and phases which are derived from the perturbative expansion, and where there exists an upper bound for the remainder, $R_{l}$, independent of the number of modes, $N$. For the Hamiltonian $H_{p}$, Eq. (1), considered in Ref. 16, the terms of the perturbative expansion could be bounded by those of a Gevrey series, which allowed to provide an exponentially small estimate for $\varepsilon^{l+1}R_{l+1}$, with $\varepsilon=A/\Delta v_{R}^{3/2}$.
@@
Note that $\Delta_{n_{i}}>1$ when $v^{\prime}=v_{0}/\Delta v_{R}$. Since we are only interested in statistical properties then, just like in Ref. 16, we estimate the remainder by its mean value with regards to the phases $\varphi_{n}$. When $v^{\prime}\approx v_{0}/\Delta v_{R}$,
@@
where $x_{\min}=\min(l)$ and $x_{\max}=\max(l)$ for each $l$ fulfilling Eq. (25). Clearly, the property of locality is satisfied if one can find an upper bound for the remainder which is independent of the extent, in phase velocity, of the wave spectrum. This is true if the integral in the right hand side of Eq. (26) converges when $x_{\max}\to\infty$. For a regular enough variation of $A^{2}(x)$ [for example when $A^{2}(x)/x$ is a monotonous function of $x$], this happens only when $A^{2}(x)/x$ goes to zero when $x$ goes to infinity, i.e., $A^{2}(v_{\phi})/v_{\phi}$ goes to zero when $v_{\phi}$ goes to infinity. Hence, the property of locality is not always true, depending on the wave spectrum.
@@
$\Delta v_{R}=\alpha\left(\overline{A^{2}_{v}}\right)^{1/3},$} (28)
with $\alpha\gg 1$, which lets naturally choose $\varepsilon=\sqrt{\overline{A^{2}_{v}}}/\Delta v_{R}^{3/2}$ for the small parameter of the perturbation analysis. This is a plain generalization of the small parameter introduced in Ref. 16 for the Hamiltonian $H_{p}$.
@@ III.2.1 The quasilinear diffusion coefficient for a discrete wave spectrum
$\Delta v_{R}=\alpha\left(D_{\rm{QL}}/\overline{k}\right)^{1/3},$ (36)
and is reminiscent of the results previously obtained by Dupree [20], although derived in a completely different sprit. Indeed, Dupree assumes diffusion in velocity in order to derive locality, which is not done here. Moreover, as discussed in Section IV, there may be diffusion even when wave-particle interaction is not local. Now,
@@
$D_{\rm{QL}}=\frac{\pi}{2}\left|\frac{k_{n}A^{2}a_{n}^{2}}{\delta v_{\phi_{n}}}\right|_{v_{\phi_{n}}\approx v},$ (39)
which yields $D_{\rm{QL}}=\pi A^{2}/2$ for the Hamiltonian dynamics defined by $H_{p}$.
@@
which is the same formula as that already published in Ref. 22.
@@ III.2.2 Validity of quasilinear diffusion and connection with locality
In order to derive quasilinear diffusion in Paragraph III.2.1, we had to assume that the dominant contribution to $\langle\Delta v^{2}\rangle$ was from the terms $n$ such that $v_{\phi_{n}}\approx v_{0}$, and that these terms all had about the same amplitude. This may only be true if the series $\sum_{n}\left[k_{n}a_{n}/(k_{n}v_{0}-\omega_{n})\right]^{2}$ converges. This is exactly the same condition as Eq. (41) for locality. Consequently, the initial quasilinear regime of diffusion (which occurs after a time interval of the order of the inverse of the total frequency spread of the wave spectrum [21]) is only expected when wave-particle interaction is local.
@@ IV.1.1 $A=1$, $v_{0}=50$
Figure 4: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=-1$, $A=1$ and $v_{0}=50$. The blue solid line is for $N=142$, the green dashed line for $N=448$, the red dashed line for $N=1415$ and the black dashed-dotted line for $N=14143$. These values for NN have been chosen so that the maximum phase velocity be, respectively, 2, 20, 200 and $2\times 10^{4}$ times $v_{0}$.
@@ IV.2.1 $A=1$, $v_{0}=50$
Figure 7: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=-2$, $A=1$ and $v_{0}=50$. The red dashed solid line is for $N=165$, the black solid line for $N=1041$. These values for NN correspond to a maximum phase velocity being, respectively, 2 and 200 times $v_{0}$.
@@
Figure 8: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=-2$, $A=11.5$ and $v_{0}=200$. The red dashed solid line is for $N=660$, the black solid line for $N=4163$. These values for NN correspond to a maximum phase velocity being, respectively, 2 and 200 times $v_{0}$.